# supplementary

### v3

@@ Abstract.
Epidemiological investigations of regionally aggregated spatial data often involve detecting spatial health disparities among neighboring regions on a map of disease mortality or incidence rates. Analyzing such data introduces spatial dependence among the health outcomes and seeks to report statistically significant spatial disparities by delineating boundaries that separate neighboring regions with disparate health outcomes. However, there are statistical challenges to appropriately defining what constitutes a spatial disparity and to construct robust probabilistic inference for spatial disparities. We enrich the familiar Bayesian linear regression framework to introduce spatial autoregression and offer model-based detection of spatial disparities. We derive exploitable analytical tractability that considerably accelerates computation. Simulation experiments conducted over a county map of the entire United States demonstrate the effectiveness of our method and we apply our method to a data set from the Institute of Health Metrics and Evaluation (IHME) on age-standardized US county-level estimates of lung cancer mortality rates.
@@ 1. Introduction
We develop statistical methods for detecting spatial disparities on maps. Disparities in health outcomes is a rapidly evolving field within public health (see, e.g., rao2023) and spatial variation in health outcomes affects how disparities manifest on disease maps. Spatial data analysis in epidemiological investigations is extensively documented (see, e.g., waller2004applied; waller2010handbook; lawson2013statistical; lawson2016handbook, and references therein). Inference on disparities proceeds from “difference boundaries” that represent significant differences among neighbors and delineate regions with significantly different disease mortality and incidence rates, which assists in improved geographic allocation of health care resources. In geographic analysis, the problem is often called “wombling” (see jacquez2003local; jacquez2003geographic; lu2005bayesian; li2011mining; ma2007bayesian; fitzpatrick2010, for some algorithmic approaches with applications). Probabilistic model-based approaches include developments in lu2007bayesian; ma2010hierarchical; li2015bayesian; banerjee2012bayesian; hanson2015spatial; leeBoundaryDetectionDisease2012; corpas-burgos2020serra; gaoSpatialDifferenceBoundary2023 and aiello2023detecting.
@@
While the aforementioned articles present an impressive range of methods, they are broadly classified into two groups. The first embraces hierarchical models and Bayesian inference for the spatial adjacency matrix by modeling its elements. The second endows the spatial random effects with a probability law that accommodates nonzero probabilities on neighboring spatial random effects being equal (often using an areal Dirichlet process or some adaptation thereof), which are computed as posterior probabilities. Both groups introduce spatial dependence and aim to balance underlying spatial similarities with the ability to detect differences among neighbors. Our current contribution offers a computationally scalable and analytically tractable approach to this problem. Instead of modeling spatial effects as realizations of a Dirichlet process, which struggles to learn from the data and impedes convergence of iterative estimation algorithms, we devise fully model-based Bayesian inference within an augmented spatial linear regression framework. More specifically, we treat spatial disparities as a multiple comparison problem within a Bayesian inferential paradigm by testing for differences in spatial random effects between neighboring regions. Subsequently, we use posterior probabilities of such differences to detect spatial disparities while controlling a Bayesian false discovery rate or FDR (mullerFDRBayesianMultiple2006). Benefits of our approach include clearer interpretation of spatial effects, computational simplicity, accounting for small differences in the values of neighboring spatial effects of adjacent regions in ascertaining spatial disparity, and delivering improved sensitivity and specificity for difference boundary detection (as seen in the analysis by gaoSpatialDifferenceBoundary2023).
@@
We exploit closed form distribution theory to account for the differences between the values of the spatial effects and evaluate posterior probabilities of such differences exceeding a certain threshold. We extend lu2005bayesian and fitzpatrick2010, who use only posterior means or spatial residuals to detect differences, to fully model-based inference. We provide an outline for the rest of the article. Section 2 discusses the problem of detecting differences between fixed effects in a linear model before we extend to spatial random effects. Section 3 develops the notion of a difference boundary, demonstrates a meaningful connection to traditional hypothesis testing in the simple linear regression model, and considers extensions to a generalized linear model. Section LABEL:sec:simulation_experiments evaluates our methods using simulated data over a map of U.S. counties followed by an analysis of lung cancer mortality data in Section LABEL:sec:application. Section LABEL:sec:_discussion concludes with a brief discussion.
@@ 2.1. Random Effects under an Augmented Linear System
We follow rieblerIntuitiveBayesianSpatial2016 who adapted the “BYM” model (besagBayesianImageRestoration1991) to ensure identifiable parameters and afford better interpretation. We write this “BYM2” model as
@@
where $\phi$ is an $n\times 1$ vector of random effects, $V_{\phi}$ is a known positive definite matrix. For subsequent analysis, we specify a conditionally autoregressive (CAR) structure, $V^{-1}_{\phi}=c(D_{W}-\alpha W)$, where $W$ is the adjacency matrix, $D_{W}$ is diagonal with the diagonal elements $D_{W,ii}$ being the number of neighbors of region $i$, $\alpha$ is the spatial smoothing parameter, and $c$ is a scaling factor. The parameter $\sigma^{2}$ captures the overall error variance if $c$ is set so that the geometric mean of the marginal prior variances is approximately one, and $\rho$ represents the proportion of total variance attributed to the spatial structure (rieblerIntuitiveBayesianSpatial2016).
@@
where $M_{\star}=(X_{\star}^{\mathrm{\scriptscriptstyle T}}V_{y_{\star}}^{-1}X_{\star})^{-1}$, $m_{\star}=X_{\star}^{\mathrm{\scriptscriptstyle T}}V_{y_{\star}}^{-1}y_{\star}=\left(\frac{1}{1-\rho}y^{\mathrm{\scriptscriptstyle T}}X+m_{0}^{\mathrm{\scriptscriptstyle T}},\frac{1}{1-\rho}y^{\mathrm{\scriptscriptstyle T}}\right)^{\mathrm{\scriptscriptstyle T}}$, $a_{n}=a_{\sigma^{2}}+\frac{n}{2}$, and $b_{n}=b_{\sigma^{2}}+\frac{1}{2}\left(\frac{1}{1-\rho}y^{\mathrm{\scriptscriptstyle T}}y+m_{0}^{\mathrm{\scriptscriptstyle T}}M_{0}m_{0}-m_{\star}^{\mathrm{\scriptscriptstyle T}}M_{\star}m_{\star}\right)$. The posterior mean $M_{\star}m_{\star}=\widehat{\gamma}_{\star}$ is the general least square regression estimate solving $X_{\star}^{\mathrm{\scriptscriptstyle T}}V_{y_{\star}}^{-1}X_{\star}\widehat{\gamma}_{\star}=X_{\star}^{\mathrm{\scriptscriptstyle T}}V_{y_{\star}}^{-1}y_{\star}$ and $b_{n}=b_{\sigma^{2}}+\frac{ns^{2}}{2}$, where $s^{2}=\frac{1}{n}(y_{\star}-X_{\star}\widehat{\gamma}_{\star})^{\mathrm{\scriptscriptstyle T}}V_{y_{\star}}^{-1}(y_{\star}-X_{\star}\widehat{\gamma}_{\star})$ is the mean residual sums of squares from (2).
@@
where $A=X^{\mathrm{\scriptscriptstyle T}}X+(1-\rho)M_{0}^{-1}$. A flat prior on $\beta$, i.e., $M_{0}^{-1}=O$, yields $A=X^{\mathrm{\scriptscriptstyle T}}X$ and
@@
These expressions are used in Section 3.2 to formulate difference boundary detection.
@@ 2.2. Limiting cases
The limits of the conditional posterior distribution of $(\beta,\gamma)$ as $\rho\rightarrow 0^{+}$ or as $\rho\rightarrow 1^{-}$ are analytically derived when the prior is non-informative on $\beta$. By the spectral decomposition, $V^{1/2}_{\phi}(I-H)V^{1/2}_{\phi}=CDC^{\mathrm{\scriptscriptstyle T}}$ where $C$ is an orthogonal matrix and $D=\text{diag}(d_{1},\ldots,d_{n})$ with entries that are the non-negative eigenvalues of $V_{\phi}^{1/2}\left(I_{n}-H\right)V_{\phi}^{1/2}$. Let $U=V^{1/2}_{\phi}C$, then $U$ is invertible, $U^{\mathrm{\scriptscriptstyle T}}V^{-1}_{\phi}U=I_{n}$, $U^{\mathrm{\scriptscriptstyle T}}(I_{n}-H)U=D$, $\operatorname{Var}(\gamma\,|\,y,\sigma^{2},\rho)=\sigma^{2}U\left(\frac{1}{\rho}I_{n}+\frac{1}{1-\rho}D\right)^{-1}U^{\mathrm{\scriptscriptstyle T}}$, and $\mathbb{E}[\gamma\,|\,y,\sigma^{2},\rho]=U\left(\frac{1-\rho}{\rho}I_{n}+D\right)^{-1}DU^{-1}y$. Calculations reveal $\lim_{\rho\rightarrow 0^{+}}\operatorname{Var}(\gamma\,|\,y,\sigma^{2},\rho)=0_{n\times n}$ and $\lim_{\rho\rightarrow 0^{+}}\mathbb{E}[\gamma\,|\,y,\sigma^{2},\rho]=0_{n}$. Therefore, if the prior on $\beta$ is flat, then $\lim_{\rho\rightarrow 0^{+}}\pi(\beta,\gamma,\sigma^{2}\,|\,y,\rho)$ collapses to the degenerate mass at $\gamma=0_{n}$.
We also obtain $\lim_{\rho\rightarrow 1^{-}}\left(\frac{1}{\rho}I_{n}+\frac{1}{1-\rho}D\right)^{-1}=I_{n}-D^{\star}$ and $\lim_{\rho\rightarrow 1^{-}}\left(\frac{1-\rho}{\rho}I_{n}+D\right)^{-1}D=D^{\star}$ where $D^{\star}\in\mathbb{R}^{n\times n}$ is diagonal with $D_{ii}^{\star}=I(d_{i}>0)$ for $i=1,\ldots,n$. Further simplifications reveal that $\mathbb{E}[\beta\,|\,y,\sigma^{2},\rho]\rightarrow(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}\left(y-UD^{\star}U^{-1}e\right)$ and $\mathbb{E}[\gamma\,|\,y,\sigma^{2},\rho]\rightarrow UD^{\star}U^{-1}e$ as $\rho\rightarrow 1^{-}$, where $e=(I-H)y$. Furthermore, $\operatorname{Var}(\beta\,|\,y,\sigma^{2},\rho)\rightarrow\sigma^{2}(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}VX(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}$, $\operatorname{Var}(\gamma\,|\,y,\sigma^{2},\rho)\rightarrow\sigma^{2}V$ and $\operatorname{Cov}(\beta,\gamma\,|\,y,\sigma^{2},\rho)\rightarrow-\sigma^{2}(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}V$ as $\rho\rightarrow 1^{-}$, where $V=V_{\phi}-UD^{\star}U^{\mathrm{\scriptscriptstyle T}}$. Finally, $\lim_{\rho\rightarrow 1^{-}}b_{n}=b_{\sigma^{2}}+\frac{1}{2}v^{\mathrm{\scriptscriptstyle T}}D^{\star}v$, where $v=U^{-1}e$ (see Section LABEL:sec:DIC_analysis). Therefore, $\lim_{\rho\rightarrow 1^{-}}\pi(\sigma^{2}\,|\,y,\rho)=\text{IG}(\sigma^{2}\,|\,a_{\sigma^{2}}+n/2,b_{\sigma^{2}}+v^{\mathrm{\scriptscriptstyle T}}D^{\star}v/2)$. These limiting cases have implications for subsequent boundary detection (see Section LABEL:subsec:epsilon_intro).

### v4

# Assessing Spatial Disparities: A Bayesian Linear Regression Approach

@@ Abstract.
Epidemiological investigations of regionally aggregated spatial data often involve detecting spatial health disparities among neighboring regions on a map of disease mortality or incidence rates. Analyzing such data introduces spatial dependence among health outcomes and seeks to report statistically significant spatial disparities by delineating boundaries that separate neighboring regions with disparate health outcomes. However, there are statistical challenges to appropriately define what constitutes a spatial disparity and to construct robust probabilistic inferences for spatial disparities. We enrich the familiar Bayesian linear regression framework to introduce spatial autoregression and offer model-based detection of spatial disparities. We derive exploitable analytical tractability that considerably accelerates computation. Simulation experiments conducted on a county map of the entire United States demonstrate the effectiveness of our method, and we apply our method to a data set from the Institute of Health Metrics and Evaluation (IHME) on age-standardized US county-level estimates of lung cancer mortality rates.
@@ 1. Introduction
We develop statistical methods for detecting spatial disparities on maps. Disparities in health outcomes are a rapidly evolving field within public health (see, e.g., rao2023) and spatial variation in health outcomes affects how disparities manifest on disease maps. Spatial data analysis in epidemiological investigations is extensively documented (see, e.g., waller2004applied; waller2010handbook; lawson2013statistical; lawson2016handbook, and references therein). Inference on disparities proceeds from “difference boundaries” that represent significant differences between neighbors and delineate regions with significantly different disease mortality and incidence rates, helping to improve the geographic allocation of health care resources. In geographic analysis, the problem is often called “wombling” (see jacquez2003local; jacquez2003geographic; lu2005bayesian; li2011mining; ma2007bayesian; fitzpatrick2010, for some algorithmic approaches with applications). Probabilistic model-based approaches include developments in lu2007bayesian; ma2010hierarchical; li2015bayesian; banerjee2012bayesian; hanson2015spatial; leeBoundaryDetectionDisease2012; corpas-burgos2020serra; gaoSpatialDifferenceBoundary2023 and aiello2023detecting.
@@
The aforementioned articles present an impressive range of methods, but are broadly classified into two groups. The first embraces hierarchical models and Bayesian inference for the spatial adjacency matrix by modeling its elements. The second endows the spatial random effects with a probability law that accommodates nonzero probabilities on neighboring spatial random effects being equal (often using an areal Dirichlet process or some adaptation thereof), which are computed as posterior probabilities. Both groups introduce spatial dependence and aim to balance the underlying spatial similarities with the ability to detect differences among neighbors. Our current contribution offers a computationally scalable and analytically tractable approach to this problem. Instead of modeling spatial effects as realizations of a Dirichlet process, which struggles to learn from the data and impedes convergence of iterative estimation algorithms, we devise fully model-based Bayesian inference within an augmented spatial linear regression framework. More specifically, we treat spatial disparities as a multiple comparison problem within a Bayesian inferential paradigm by testing for differences in spatial random effects between neighboring regions. Subsequently, we use posterior probabilities of such differences to detect spatial disparities while controlling a Bayesian false discovery rate or FDR (mullerFDRBayesianMultiple2006). The benefits of our approach include clearer interpretation of spatial effects, computational simplicity, accounting for small differences in the values of neighboring spatial effects of adjacent regions in determining spatial disparity, and providing improved sensitivity and specificity for the detection of boundary differences (as seen in the analysis by gaoSpatialDifferenceBoundary2023). Our parametric approach is suitable for applications where one disease map is of interest and lack of replication hinders non-parametric methods (such as in slackGeographicConcentrationUs2014; jagai2017county; rodriguez2018county).
@@
We exploit closed-form distribution theory to account for the differences between the values of the spatial effects and evaluate posterior probabilities of such differences exceeding a certain threshold. We extend lu2005bayesian and fitzpatrick2010, who use only posterior means or spatial residuals to detect differences, to fully model-based inference. We provide an outline for the rest of the article. Section 2 discusses the problem of detecting differences between fixed effects in a linear model before we extend to spatial random effects. Section 3 develops the notion of a difference boundary, demonstrates a meaningful connection to traditional hypothesis testing in the simple linear regression model, and considers extensions to a generalized linear model. Section LABEL:sec:simulation_experiments evaluates our methods alongside an existing Bayesian nonparametric approach using simulated data over a map of Californian counties followed by an analysis of lung cancer mortality data in Section LABEL:sec:application. Section LABEL:sec:_discussion concludes with a brief discussion.
@@ 2.1. Random Effects under an Augmented Linear System
We follow rieblerIntuitiveBayesianSpatial2016 who adapted the “BYM” model (besagBayesianImageRestoration1991) to ensure identifiable parameters and provide a better interpretation. We write this “BYM2” model as
@@
where $\phi$ is an $n\times 1$ vector of random effects, $V_{\phi}$ is a known positive definite matrix such that $\pi(\phi)$ is proper. The parameter $\sigma^{2}$ captures the overall error variance if $c$ is set so that the geometric mean of the marginal prior variances is equal to one, and $\rho$ represents the proportion of total variance attributed to the spatial structure (rieblerIntuitiveBayesianSpatial2016). For subsequent analysis, we specify a conditionally autoregressive (CAR) structure, one of the most popular spatial models used in areal data analysis. Here, $\pi(\gamma_{i}\,|\,\sigma^{2},\gamma_{j},j\neq i)=N(\alpha\sum_{j}w_{ij}\gamma_{j}/w_{i+},c\sigma^{2}/w_{i+})$, where $w_{ij}$ is the binary adjacency indicator for regions $i$ and $j$, $\alpha$ is a spatial smoothing parameter, $c$ is a fixed scaling constant and $w_{i+}=\sum_{j}w_{ij}$ is the number of neighbors of region $i$. The conditional distribution of each spatial residual is based on a linear combination of its neighbors. By Brook’s lemma (besagBayesianImageRestoration1991), this equates to setting $V^{-1}_{\phi}=c(D_{W}-\alpha W)$ in (1), where $W$ is the adjacency matrix, $D_{W}$ is diagonal with the diagonal elements $D_{W,ii}=w_{i+}$ (banerjeeHierarchicalModelingAnalysis2015, Chap. 4). The density $\pi(\phi)$ is proper when $\alpha\in(1/\lambda_{min},1)$, where $\lambda_{min}$ is the smallest eigenvalue of $D_{W}^{-1/2}WD_{W}^{-1/2}$. Since $\rho$ already controls the amount of spatial smoothing in (1), we set $\alpha=0.99$ in the following sections to improve the identifiability of spatial effects $\phi$, which we consider as potential indicators of latent factors driving local health inequity.
@@
where $M_{\star}=(X_{\star}^{\mathrm{\scriptscriptstyle T}}V_{y_{\star}}^{-1}X_{\star})^{-1}$, $m_{\star}=X_{\star}^{\mathrm{\scriptscriptstyle T}}V_{y_{\star}}^{-1}y_{\star}=\left(\frac{1}{1-\rho}y^{\mathrm{\scriptscriptstyle T}}X+m_{0}^{\mathrm{\scriptscriptstyle T}},\frac{1}{1-\rho}y^{\mathrm{\scriptscriptstyle T}}\right)^{\mathrm{\scriptscriptstyle T}}$, $a_{n}=a_{\sigma^{2}}+\frac{n}{2}$, and $b_{n}=b_{\sigma^{2}}+\frac{1}{2}\left(\frac{1}{1-\rho}y^{\mathrm{\scriptscriptstyle T}}y+m_{0}^{\mathrm{\scriptscriptstyle T}}M_{0}m_{0}-m_{\star}^{\mathrm{\scriptscriptstyle T}}M_{\star}m_{\star}\right)$. The posterior mean $M_{\star}m_{\star}=\widehat{\gamma}_{\star}$ is the general least squares regression estimate to be solved $X_{\star}^{\mathrm{\scriptscriptstyle T}}V_{y_{\star}}^{-1}X_{\star}\widehat{\gamma}_{\star}=X_{\star}^{\mathrm{\scriptscriptstyle T}}V_{y_{\star}}^{-1}y_{\star}$ and $b_{n}=b_{\sigma^{2}}+\frac{ns^{2}}{2}$, where $s^{2}=\frac{1}{n}(y_{\star}-X_{\star}\widehat{\gamma}_{\star})^{\mathrm{\scriptscriptstyle T}}V_{y_{\star}}^{-1}(y_{\star}-X_{\star}\widehat{\gamma}_{\star})$ is the mean residual sums of squares from (2).
@@
where $A=X^{\mathrm{\scriptscriptstyle T}}X+(1-\rho)M_{0}^{-1}$. A flat prior to $\beta$, that is, $M_{0}^{-1}=O$, yields $A=X^{\mathrm{\scriptscriptstyle T}}X$ and
@@
These expressions are used in Section 3.2 to formulate the detection of differences between boundaries.
@@ 2.2. Limiting cases
The limits of the conditional posterior distribution of $(\beta,\gamma)$ as $\rho\rightarrow 0^{+}$ or as $\rho\rightarrow 1^{-}$ are analytically derived when the prior is non-informative on $\beta$. By spectral decomposition, $V^{1/2}_{\phi}(I-H)V^{1/2}_{\phi}=CDC^{\mathrm{\scriptscriptstyle T}}$ where $C$ is an orthogonal matrix and $D=\text{diag}(d_{1},\ldots,d_{n})$ with entries that are nonnegative eigenvalues of $V_{\phi}^{1/2}\left(I_{n}-H\right)V_{\phi}^{1/2}$. Let $U=V^{1/2}_{\phi}C$, then $U$ is invertible, $U^{\mathrm{\scriptscriptstyle T}}V^{-1}_{\phi}U=I_{n}$, $U^{\mathrm{\scriptscriptstyle T}}(I_{n}-H)U=D$, $\operatorname{Var}(\gamma\,|\,y,\sigma^{2},\rho)=\sigma^{2}U\left(\frac{1}{\rho}I_{n}+\frac{1}{1-\rho}D\right)^{-1}U^{\mathrm{\scriptscriptstyle T}}$, and $\mathbb{E}[\gamma\,|\,y,\sigma^{2},\rho]=U\left(\frac{1-\rho}{\rho}I_{n}+D\right)^{-1}DU^{-1}y$. Calculations reveal $\lim_{\rho\rightarrow 0^{+}}\operatorname{Var}(\gamma\,|\,y,\sigma^{2},\rho)=0_{n\times n}$ and $\lim_{\rho\rightarrow 0^{+}}\mathbb{E}[\gamma\,|\,y,\sigma^{2},\rho]=0_{n}$. Therefore, if the prior on $\beta$ is flat, then $\lim_{\rho\rightarrow 0^{+}}\pi(\beta,\gamma,\sigma^{2}\,|\,y,\rho)$ collapses to the degenerate mass at $\gamma=0_{n}$.
We also obtain $\lim_{\rho\rightarrow 1^{-}}\left(\frac{1}{\rho}I_{n}+\frac{1}{1-\rho}D\right)^{-1}=I_{n}-D^{\star}$ and $\lim_{\rho\rightarrow 1^{-}}\left(\frac{1-\rho}{\rho}I_{n}+D\right)^{-1}D=D^{\star}$ where $D^{\star}\in\mathbb{R}^{n\times n}$ is diagonal with the entries $D_{ii}^{\star}=I(d_{i}>0)$ for $i=1,\ldots,n$. Further simplifications reveal that $\mathbb{E}[\beta\,|\,y,\sigma^{2},\rho]\rightarrow(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}\left(y-UD^{\star}U^{-1}e\right)$ and $\mathbb{E}[\gamma\,|\,y,\sigma^{2},\rho]\rightarrow UD^{\star}U^{-1}e$ as $\rho\rightarrow 1^{-}$, where $e=(I-H)y$ are the residuals of ordinary least squares. Furthermore, $\operatorname{Var}(\beta\,|\,y,\sigma^{2},\rho)\rightarrow\sigma^{2}(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}VX(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}$, $\operatorname{Var}(\gamma\,|\,y,\sigma^{2},\rho)\rightarrow\sigma^{2}V$ and $\operatorname{Cov}(\beta,\gamma\,|\,y,\sigma^{2},\rho)\rightarrow-\sigma^{2}(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}V$ as $\rho\rightarrow 1^{-}$, where $V=V_{\phi}-UD^{\star}U^{\mathrm{\scriptscriptstyle T}}$. Finally, $\lim_{\rho\rightarrow 1^{-}}b_{n}=b_{\sigma^{2}}+\frac{1}{2}v^{\mathrm{\scriptscriptstyle T}}D^{\star}v$, where $v=U^{-1}e$. Therefore, $\lim_{\rho\rightarrow 1^{-}}\pi(\sigma^{2}\,|\,y,\rho)=\text{IG}(\sigma^{2}\,|\,a_{\sigma^{2}}+n/2,b_{\sigma^{2}}+v^{\mathrm{\scriptscriptstyle T}}D^{\star}v/2)$. These limiting cases have implications for subsequent boundary detection (see Section LABEL:subsec:epsilon_intro).